OpenStudy (anonymous):

Find all the solutions in the interval [0,2pi) for cos2x= 1/cos^2x - sin^2x

OpenStudy (anonymous):

\[\cos2x = 1\div \cos ^{2}x-\sin ^{2}x\] here is the equation again

OpenStudy (anonymous):

does it help to know that \[\cos(2x)=\cos^2(x)-\sin^2(x)\]?

OpenStudy (anonymous):

i know that its the double angle formula

OpenStudy (anonymous):

@bridgetx516x \[\cos 2x=\cos^2 x-\sin^2 x\]

OpenStudy (anonymous):

so you have \[\cos^2(x)-\sin^2(x)=\frac{1}{\cos^2(x)-\sin^2(x)}\]

OpenStudy (anonymous):

yes but i didnt know if i should bring it all to one side or multiply by the denominator

OpenStudy (anonymous):

the only way for a number to be equal to its reciprocal is if it is \(1\) or \(-1\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so instead of multiplying a bunch of stuff out, you know \[\cos^2(x)-\sin^2(x)=1\] or \[\cos^2(x)-\sin^2(x)=-1\]

OpenStudy (anonymous):

oh i wasnt sure if cos2(x)−sin2(x)=−1 but that makes sense now

OpenStudy (anonymous):

so either sine is zero and cosine is one, or cosine is zero and sine is one

OpenStudy (anonymous):

i think cosine is zero and sine is one

OpenStudy (anonymous):

sure for example at \(\frac{\pi}{2}\) you have \[\cos^2(\frac{\pi}{2})-\sin^2(\frac{\pi}{2})=-1\]

OpenStudy (anonymous):

but also at \(0\) you have \[\cos^2(0)-\sin^2(0)=1-0=1\] so that one works as well

OpenStudy (anonymous):

so would my final answer be 0, pi/2, pi, 3pi/2

OpenStudy (anonymous):

\[\cos^2 2x=1\] which means cos 2x=\[\pm1\] in[0,2pi) there are 2 solution one 0 and another pi

OpenStudy (anonymous):

also pi/2 and 3*pi/2

OpenStudy (anonymous):

4 solutns whole

OpenStudy (anonymous):

yes for solutions, you have them you can check that they work

OpenStudy (phi):

btw, even if you multiplied both sides by cos^2 - sin^2 \[ \cos^2(x)-\sin^2(x)=\frac{1}{\cos^2(x)-\sin^2(x)} \] you would get \[ (\cos^2(x)-\sin^2(x))^2 = 1\] take the square root, and you get \[ \cos^2(x)-\sin^2(x) = ±1\] just as satellite said you would at this point , if you replaced cos^2 with 1- sin^2 you can solve for sin(x) if you did not notice that 0, and ±1 for sin and cos are the solutions

OpenStudy (anonymous):

than you everyone!